OpenStudy (anonymous):

A galvanized iron sheet 64 cm wide is to be bent to form a gutter by turning up vertical strips along each side. How many centimeters should be turned up at each side in order to give a gutter with the maximum cross-sectional area?

OpenStudy (darkbluechocobo):

Your teacher must really hate giving out details for problems ._.

OpenStudy (anonymous):

Yep :) It really gives me a headache >.<

OpenStudy (darkbluechocobo):

I could imagine these are table flipping problems

OpenStudy (darkbluechocobo):

Which luckily I can help you :p

OpenStudy (darkbluechocobo):

Two dimensions: v = the vertical height of the turned up edge b = the width of the remaining base. 2*v + b = 64 v * b = A (objective Amax) Let's simplify the second equation by substituting b from the 1st: b = 64 - 2*v, so v * (64 - 2*v) = A, or 64*v - 2*v^2 = A(v) We want to know the maximum A(v), which can be found where the derivative of A(v) with respect to v is 0: d[A(v)] = 64 - 4*v, where 64 - 4*v = 0, v = 16 Test this: using equation 1, 2*v + b = 64, or 2 * 16 +b = 64, b = 32 So Amax would be v * b, or 16 * 32 = 512 cm^2, (whereas dividing the sheet in 3rds, giving height and width of 21.33 cm will only provide 455 cm^2)

OpenStudy (darkbluechocobo):

455 cm^2<<< Basically the answer

OpenStudy (darkbluechocobo):

Basically the short cut to reading all dem words

OpenStudy (anonymous):

Thank you so much! @DarkBlueChocobo :)

OpenStudy (darkbluechocobo):

Anything I can help you with I shall try :D Even though I must say I want to flip at least 12 tables at your teacher >.<

OpenStudy (anonymous):

Hahaha xD I would like to do that too but I can't >.< @DarkBlueChocobo

OpenStudy (darkbluechocobo):

*sigh* Well someone needs too I can contact my table flipping friend if you need me too XD

OpenStudy (anonymous):

Really? Haha xD Who?

OpenStudy (darkbluechocobo):

Well my friend Eliot and I can teach him a lesson. Dem Table Flips hurt people. ._. Just tie him up in the middle of the night and put him in a room full of tables. That is my dream.

OpenStudy (darkbluechocobo):

Dem Tables be hurting

OpenStudy (darkbluechocobo):

Not to mention those tables would be made out of MAHOGANY

OpenStudy (anonymous):

It would really hurt

OpenStudy (darkbluechocobo):

Me and my friend would be like ┻━┻ ︵ヽ(`Д´)ノ︵ ┻━┻

OpenStudy (darkbluechocobo):

Then after that we would flip your teacher (╯°□°)╯ /(.□. \)

OpenStudy (anonymous):

Great! Really. It will the best show I will ever see! xD

OpenStudy (darkbluechocobo):

Yes :p xD o.o Sparkles (ノ◕ヮ◕)ノ*:・゚✧*:・゚✧

OpenStudy (darkbluechocobo):

Well my table flipping friend any more evil math questions you would like me to help you conquer

OpenStudy (anonymous):

Sure! No problem! :)

OpenStudy (anonymous):

It will be an honor

OpenStudy (darkbluechocobo):

Also here have some sparkles It will do you some good with a teacher like that (ノ◕ヮ◕)ノ*:・゚✧*:・゚✧

OpenStudy (anonymous):

Amazing! ^_^

OpenStudy (darkbluechocobo):

^-^ mhm the magic of sparkles

OpenStudy (anonymous):

Yes, the magic of sparkles!

OpenStudy (darkbluechocobo):

*poke* ._.

OpenStudy (anonymous):

By the way, I would like to say, thank you for helping me today. I need to go. See you! :)

OpenStudy (darkbluechocobo):

Alright :3 Bye It has been a pleasure helping you out